aboyles:

Solve the following problem. It may be helpful to draw a chart on scrap paper to organize the information and write the equation. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. A clerk was asked to change a $10 bill. She returned 9 more dimes than nickels and twenty-one more quarters than dimes. How many coins of each did she return?

GalaxyzStarz:

Let "x" = Be the number of nickels \[x + 9 = dimes \] (nine more dimes than nickels \[x + 9 + 21 = quarters \] (21 more quarters than dimes) \[ .05(x) + .1(x+9) + .25(x + 9 + 21) = 10\] \[.4x + 8.4 = 10\] \[.4x = 1.6\] \[x = 4 \rightarrow nickels\] \[x+ 9 = 13 \rightarrow dimes\] \[x + 30 = 34 \rightarrow quarters\] I hope this clears up some questions you may had.